Project

General

Profile

Actions

Commissioning and Ops Scheduling meeting : 19/1/2016

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Plans for realignment of SSD
    • Test of neutron monitor whilst we have access
    • Simulation of effect of air in the drift space between tracker and absorber (John N)
  • Upcoming and ongoing activities : chaired by DC
  • Issues for discussion
    • Cooling water controls update (John G)
    • BLOC and expert recruitment : Next user cycle is Feb 16 - Mar 25. Shifts will start on Feb 23rd - need BLOCs and SOCs.
  • AOB
  • Date of next meeting : January 26th

Minutes

Present: Paul H, Chris R, Mark T, Paolo, Steve B, Colin, Ken, Henry, Pierrick, John N, John G, Andy N, Josef, Ryan, Mike C

Review of tasks and activities identified at last meeting

  • Realignment : John G and Andy have been doing lots of preliminary work on the survey. Plans for reconstruction afterrealignment are underway. There are 3 plans to replace the bellows : an inflateable seal (which may not be ready on the timescale of the 2015/04 user cycle),
    a fully engineered solution for an offset bellows (marginal for 2015/04; drawings
    are almost ready and order could go out this week, but there is a 6-8 week delivery schedule), and retrofitting the current
    bellows. This is doable in the time available. The physics work (John Cobb) still needs to be done. We are putting the channel elements
    into an accelerator physics alignment model to validate the alignment features.

Chris : Emphasise that we need a survey report at the end of the procedure.

Action Item (Colin to prod John C) : need to get John's data with fiducials in it to get to Tony Middleton.

  • Simulation of air in the tracker (John N) :

John has looked through all existing simulations with (200, 6pi) : one with Xe in the absorber and air, He and 50/50
in the DS tracker volume. Preliminary look suggests that air in the tracker volume makes minimal impact but he will collate
the plots into a talk.

Action Item : John to give these files to the tracker group to make certain tracker resolution doesn't change
Action Item : John to do a simulation with air in the interspace and He in the absorber volume
Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
Action Item : Colin will check with Alan Bross whether the LiH absorber can withstand vacuum

  • neutron monitor test : John says that this is going to be booked in and Health Physics will look at it before the startup.

Upcoming and ongoing activities

  • The PPS will be serviced on Tuesday 26th January. This is likely to take about 8 hours and the Hall will be closed during this.
  • We need to ensure that all work in the DSA is down before February 1st. Close up will start on the 29th.
  • Linde fridge will be cooled down starting Monday 25th.

Action Item : John will send Steve his plan for the next two weeks to enter on the operations calendar.

Issues for discussion

  • Cooling water discussion - There is a plan to implement cooling water controls which will allow us to move compressors from
    current cooling circuit onto a separate one, which will be better for the compressors since they will be running with glycol
    and they'll be separate from the conventional magnets. Scheduled for February (?). However there seems to have been a miscommunication
    between Colin and John on the details of the plan.

Action Item : Colin and John to talk and agree on the plan.

  • ISIS monitoring - how long will we let ISIS monitor the magnets?

Action Item : Josef/Colin to check with Alan about monitoring requirements for the SS magnets. We should plan to remove the checks to
play fair with ISIS. Perhaps Alan will let us keep the buzzer, but remove the active monitoring.

  • BLOCs and SOCs need to be scheduled for the upcoming user cycle. Henry will try to recruit more BLOCs and do some training in the next
    few weeks

Action Item : Steve to remind Adam that we need SOCs.

AOB

  • (Colin) MICE is moving to R1 early in February. Room for 24 desks and a seperate room with a lockable door for the external control room.
    Already boxes in ATLAS and Colin has labels. MICE area is on the top of R1.
  • (Ken) : Optics review went very well. Ken was planning to send out a note with some of the key plots and the way everyone dealt
    with questions was very good. Publication will take place in 4 months.
  • (Colin) : DS power supply is still coming back on "end of February". Time for commissioning is unknown. Mike recommends that the
    full commissioning work be done from scratch. This could take up to 2 weeks. If this is the case we will need to postpone this until
    after the user cycle.

Action Item : Colin is to get an estimate for power supply installation time.

added after the meeting : Steve Griffiths wishes to do SS performance tests on 1/2 Feb. We require the Hall to be closed. Steve
should talk to the DC about scheduling.

Updated by Boyd, Steven almost 8 years ago ยท 17 revisions